# Tutorial on amortized optimization

### v4

@@ 3.6.3 The deterministic policy gradient
The policy learning of many model-free actor-critic reinforcement learning algorithms on continuous spaces can be seen as objective-based amortization of the max-$Q$ operation in eq. 3.37. This includes the policy improvement step for deterministic policies as pioneered by the deterministic policy gradient (DPG) by Silver et al. [2014], deep deterministic policy gradient (DDPG) by Lillicrap et al. [2016], and and twin delayed DDPG (TD3) by Fujimoto et al. [2018]. All of these methods interweave 1) policy evaluation to estimate the state-action value $Q^{\pi_{\theta}}$ of the current policy $\pi_{\theta}$, and 2) policy improvement to find the policy that best-optimizes the max-Q operation with
@@ 3.6.4 The stochastic value gradient and soft actor-critic
This amortization viewpoint can be extended to stochastic systems and policies that use the stochastic value gradient (SVG) and covers a range of model-free to model-based method depending on how the value is estimated [Byravan et al., 2019, Hafner et al., 2020, Byravan et al., 2022, Amos et al., 2021]. As observed in Haarnoja et al. [2018], Amos et al. [2021], the policy update step in the soft actor-critic can also be seen as a model-free stochastic value gradient with the value estimate regularized or “softened” with an entropy term. These methods learn policies $\pi_{\theta}$ that amortize the solution to a stochastic optimization problem, such as eqs. 3.39 and 3.40, over some distribution of states $p(x)$, such as the stationary state distribution or an approximation of it with a replay buffer. Taking the expected value under the policy gives the policy loss
@@
which softens the policy and value estimates with an entropy regularization as in Haarnoja et al. [2018], Amos et al. [2021]. Eq. 3.50 can be seen as an entropy-regularized value estimate by expanding the KL
@@
Mohamed et al. [2020] discusses standard ways of optimizing eqs. 3.49 and 3.50, which could be with a likelihood ratio gradient estimator [Williams, 1992] or via reparameterization.
@@ 3.6.6 Guided policy search (GPS)
The GPS family of methods [Levine and Koltun, 2013, Levine and Abbeel, 2014, Levine et al., 2016, Montgomery and Levine, 2016] fits an approximate dynamics model to data and then amortizes the solutions of a controller that solves the MPC problem in eq. 3.38 within a local region of the data to improve the amortization model. Given samples of $\pi^{\star}$ from a controller that solves eq. 3.40, GPS methods typically optimize the KL divergence between the controller and these samples with
@@ 3.6.7 POPLIN by Wang and Ba [2020]
Control. Next they explore ways of using the learned policy $\pi_{\theta^{\star}}$ to solve the model predictive control optimization problem in eq. 3.38. The POPLIN-A variant (“A” stands for “action”) uses $\pi_{\theta^{\star}}$ to predict an initial control sequence $\hat{u}^{0}_{1:H}$ that is then passed as an input to a controller based on the cross-entropy method over the action space that uses a learned model on the trajectories. The POPLIN-P variant (“P” stands for “parameter”) suggests that the parameter space of the fully-amortized model has learned useful information about the structure of the optimal action sequences. As an alternative to solving the MPC problem in eq. 3.38 over the action space, POPLIN-P proposes to use CEM to find a perturbation $\omega$ to the optimal parameters $\theta^{\star}$ that maximizes the value from a state $x$ with
@@
Thus the action produced by $\pi_{\theta^{\star}+\omega^{\star}}(x)$ is a solution the control problem in eq. 3.38 obtained by adapting the policy’s parameters to the state $x$.
@@
The initial phase is fully-amortized behavioral cloning. The second fine-tuning phase can seen as semi-amortization that learns only the initialization $\theta$ and finds $\omega$ with CEM, with a regression-based loss ${\mathcal{L}}_{\rm reg}$ that only has knowledge of the initial model and does not include the adaptation.
@@ 3.6.8 The differentiable cross-entropy method by Amos and Yarats [2020]
Amos and Yarats [2020] considers a semi-amortized learning setting that learns a latent domain for control, which can be seen as a similarly-motivated alternative to the parameter-space control done in POPLIN. The key piece of latent control is to learn a decoder $\varphi_{\theta}:{\mathcal{Z}}\rightarrow{\mathcal{U}}^{H}$ that maps from a low-dimensional latent space ${\mathcal{Z}}$ to the $H$-step control space that solves eq. 3.38. Learning a latent space is useful if there are many redundancies and possibly bad local minima on the original control space ${\mathcal{U}}^{H}$ that the latent space can get rid of. Given an initial state $x_{1}$, the optimal latent representation can be obtained by solving the control optimization problem over ${\mathcal{Z}}$ with
@@
Solving eq. 3.56 with DCEM enables the optimal solution $\hat{z}_{\theta}(x)$ to be differentiated with respect to the parameters $\theta$ of the decoder $\varphi_{\theta}$. The final predicted control sequence can be obtained with the decoder $\hat{u}_{1:H}(x;\theta)\vcentcolon=\varphi_{\theta}(\hat{z}_{\theta}(x))$ and the decoder can be learned by regressing onto ground-truth control sequences $u^{\star}(x)$ with
@@
Figure 3.3 visualizes an example on the cartpole task where this is able to learn a latent space that captures the cyclic and smoothness structure of the optimal control sequence space.
@@ 3.6.10 Learning the value function
The $Q$-value function eq. 3.36 for finding an MDP policy is often unknown and needs to be estimated from data along with the policy $\pi$ that is amortizing it, e.g. in eq. 3.37. Actor-critic methods are a common reinforcement learning approach that jointly learn a policy $\pi$ (the actor) and $Q$-value estimate (the critic) [Konda and Tsitsiklis, 1999, Sutton and Barto, 2018]. The policy amortizes the current $Q$ estimate, e.g. by using an approach previously discussed in this section, and the $Q$ function is fit to data sampled from the system. One way of learning the $Q$ function is to replace the value estimate $V^{\pi}$ in eq. 3.36 with the $Q$-value estimate to yield the relationship
@@
which is referred to as the Bellman equation [Bellman, 1966, Sutton and Barto, 2018]. Eq. 3.61 is an equality that should hold over all states and actions in the system and a value estimate can be parameterized and learned to satisfy the relationship. While the best way of learning the $Q$ estimate is an open research topic [Watkins and Dayan, 1992, Baird, 1995, Ernst et al., 2005, Maillard et al., 2010, Scherrer, 2010, Geist et al., 2017, Le et al., 2019, Fujimoto et al., 2022], a common way is to optimize residual of eq. 3.61 with
@@
VAE on MNIST (4.1.1) $784$ (=$28\cdot 28$, MNIST digits) $20$ (parameterizing a 10D Gaussian)
@@
Model-based control (4.1.3) $45$ (humanoid states) $51$ (=$17\cdot 3$, short action sequence)
Sphere (4.2) $16$ ($c$-convex function parameterizations) $3$ (sphere)
Turning now to the implementation details, this section looks at how to develop and analyze amortization software. The standard and easiest approach in most settings is to use automatic differentiation software such as Maclaurin et al. [2015a], Al-Rfou et al. [2016], Abadi et al. [2016], Bezanson et al. [2017], Agrawal et al. [2019b], Paszke et al. [2019], Bradbury et al. [2020] to parameterize and learn the amortization model. There are many open source implementations and re-implementations of the methods in chapter 3 that provide a concrete starting point to start building on them. This section looks closer at three specific implementations: section 4.1 evaluates the amortization components behind existing implementations of variational autoencoders section 3.1 and control section 3.6 and section 4.2 implements and trains an amortization model to optimize functions defined on a sphere. Table 4.1 summarizes the concrete dimensions of the amortization problems considered here and section 4.3 concludes with other useful software references. The source code behind this section is available at https://github.com/facebookresearch/amortized-optimization-tutorial.
@@ 4.1 Amortization in the wild: a deeper look
This section shows code examples of how existing implementations using amortized optimization define and optimize their models for variational autoencoders (section 4.1.1) and control and policy learning (sections 4.1.2 and 4.1.3). The amortization component in these systems is often a part of a larger system to achieve a larger task: VAEs also reconstruct the source data after amortizing the ELBO computation in eq. 3.5 and policy learning methods also estimate the $Q$-value function in section 3.6.10. This section scopes to the amortization components to show how they are implemented. I have also added evaluation code to the pre-trained amortization models from existing repositories and show that the amortized approximation often obtains a solution up to 25000 times faster than solving the optimization problems from scratch on an NVIDIA Quadro GP100 GPU.
@@ 4.1.1 The variational autoencoder (VAE)
This section looks at the code behind standard VAE [Kingma and Welling, 2014] that follows the amortized optimization setup described in section 3.1.1. While there are many implementations for training and reproducing a VAE, this section will use the implementation at https://github.com/YannDubs/disentangling-vae, which builds on the code behind Dupont [2018] at https://github.com/Schlumberger/joint-vae. While the repository is focused on disentangled representations and extensions of the original VAE formulation, this section only highlights the parts corresponding to the original VAE formulation. The code uses standard PyTorch in a minimal way that allow us to easily look at the amortization components.
Training the VAE. Figure 4.2 paraphrases the relevant snippets of code to implement the main amortization problem in eq. 3.4 for image data where the likelihood is given by a Bernoulli. Figure 4.2(a) defines an encoder $\hat{\lambda}_{\theta}$, to predicts a solution to the ELBO implemented in fig. 4.2(b), which is optimized in a loop over the training data (images) in fig. 4.2(c). The README in the repository contains instructions for running the training from scratch. The repository contains the binary of a model trained on the MNIST dataset [LeCun, 1998], which the next portion evaluates.
@@
Figure 4.1: Runtime comparison between Adam and an amortized encoder $\hat{\lambda}_{\theta}$ to solve eq. 3.3 for a VAE on MNIST. This uses a batch of 1024 samples and was run on an unloaded NVIDIA Quadro GP100 GPU. The values are normalized so that $\lambda(x)=0$ takes a value of -1 and the optimal $\lambda^{\star}$ takes a value of 0. The amortized policy is approximately 25000 times faster than solving the problem from scratch.
1class Encoder(nn.Module): # From disvae.models.encoders
@@
(b) Definition of the $\operatorname{ELBO}$ in eq. 3.1
@@
Evaluating the VAE. This section looks at how well the amortized encoder $\hat{\lambda}$ approximates the optimal encoder $\lambda^{\star}$ given by explicitly solving eq. 3.3, which is referred to the amortization gap [Cremer et al., 2018]. eq. 3.3 can be solved with a gradient-based optimizer such as SGD or Adam [Kingma and Ba, 2015]. Figure 4.4 shows the key parts of the PyTorch code for making this comparison, which can be run on the pre-trained MNIST VAE with code/evaluate_amortization_speed_function_vae.py.
Figure 4.1 shows that the amortized prediction from the VAE’s encoder predicts the solution to the ELBO 25000 times faster (!) than running 2k iterations of Adam on a batch of 1024 samples. This is significant as every training iteration of the VAE requires solving eq. 3.3, and a large model may need millions of training iterations to converge. Amortizing the solution makes the difference between the training code running in a few hours instead of a few months if the problem was solved from scratch to the same level of optimality. Knowing only the ELBO values is not sufficient to gauge the quality of approximate variational distributions. To help understand the quality of the approximate solutions, fig. 4.3 plots out the decoded samples alongside the original data.
@@
Figure 4.4: Evaluation code for comparing the amortized prediction $\hat{y}$ to the true solution $y^{\star}$ solving eq. 1.1 with a gradient-based optimizer. The full instrumented version of this code is available in the repository associated with this tutorial at code/evaluate_amortization_speed_function.py.
@@ 4.1.2 Control with a model-free value estimate
Figure 4.5: Runtime comparison between Adam and a learned policy $\pi_{\theta}$ to solve eq. 3.37 on the humanoid MDP. This was evaluated as a batch on an expert trajectory with 1000 states and was run on an unloaded NVIDIA Quadro GP100 GPU. The values are normalized so that $\pi(x)=0$ takes a value of -1 and the optimal $\pi^{\star}$ takes a value of 0. The amortized policy is approximately 1000 times faster than solving the problem from scratch.
@@
This section dives into the training and evaluation code for learning a deterministic model-free policy $\pi_{\theta}:{\mathcal{X}}\rightarrow{\mathcal{Y}}$ to amortize a model-free value estimate $Q$ for controlling the humanoid MDP from Brockman et al. [2016] visualized in fig. 1.2. This MDP has $|{\mathcal{X}}|=45$ states (angular positions and velocities describing the state of the system) and $|{\mathcal{Y}}|=17$ actions (torques to apply to the joints). A model-free policy $\pi$ maps the state to the optimal actions that maximize the value on the system. Given a known action-conditional value estimate $Q(x,u)$, the optimal policy $\pi^{\star}$ solves the optimization problem in eq. 3.37 that the learned policy $\pi$ tries to match, e.g. using policy gradient in eq. 3.47.
@@
The codebase behind Amos et al. [2021] at https://github.com/facebookresearch/svg contains trained model-free policy and value estimates on the humanoid in addition to model-based components the next section will use. The full training code there involves parameterizing a stochastic policy and estimating many additional model-based components, but the basic training loop for amortizing a deterministic policy from the solution there can be distilled into a form similar to fig. 4.2.
@@
This section mostly focuses on evaluating the performance of the trained model-free policy in comparison to maximizing the model-free value estimate eq. 3.37 from scratch for every state encountered. An exhaustive evaluation of a solver for eq. 3.37 would need to ensure that the solution is not overly adapted to a bad part of the $Q$ estimate space — because $Q$ is also a neural network susceptible to adversarial examples, it is very likely that directly optimizing eq. 3.37 may result in a deceptively good policy when looking at the $Q$ estimate that does not work well on the real system. For simplicity, this section ignores these issues and normalizes the values to $[-1,0]$ where $-1$ will correspond to the value from taking a zero action and 00 will correspond to the value from taking the expert’s action. (This is valid in this example because the zero action and expert action never coincide.)
@@
Figure 4.5 shows that the amortized policy is approximately 1000 times faster than solving the problem from scratch. The $Q$ values presented there are normalized and clamped so that the expert policy has a value of zero and the zero action has a value of -1. This example can be run with code/evaluate_amortization_speed_function_control.py, which shares the evaluation code also used for the VAE in fig. 4.4.

### v5

@@ 3.6.3 The deterministic policy gradient
The policy learning of many model-free actor-critic reinforcement learning algorithms on continuous spaces can be seen as objective-based amortization of the max-$Q$ operation in eq.˜3.37. This includes the policy improvement step for deterministic policies as pioneered by the deterministic policy gradient (DPG) by Silver et al. [2014], deep deterministic policy gradient (DDPG) by Lillicrap et al. [2016], and twin delayed DDPG (TD3) by Fujimoto et al. [2018]. All of these methods interweave 1) policy evaluation to estimate the state-action value $Q^{\pi_{\theta}}$ of the current policy $\pi_{\theta}$, and 2) policy improvement to find the policy that best-optimizes the max-Q operation with
@@ 3.6.4 The stochastic value gradient and soft actor-critic
This amortization viewpoint can be extended to stochastic systems and policies that use the stochastic value gradient (SVG) and covers a range of model-free to model-based method depending on how the value is estimated [Byravan et al., 2019, Hafner et al., 2020, Byravan et al., 2022, Amos et al., 2021]. As observed in Haarnoja et al. [2018], Amos et al. [2021], the policy update step in the soft actor-critic can also be seen as a model-free stochastic value gradient with the value estimate regularized or “softened” with an entropy term. These methods learn policies $\pi_{\theta}$ that amortize the solution to a stochastic optimization problem, such as eqs.˜3.39 and 3.40, over some distribution of states $p(x)$, such as the stationary state distribution or an approximation of it with a replay buffer. Taking the expected value under the policy gives the policy loss
@@
which softens the policy and value estimates with an entropy regularization as in Haarnoja et al. [2018], Amos et al. [2021]. Eq.˜3.50 can be seen as an entropy-regularized value estimate by expanding the KL
@@
Mohamed et al. [2020] discusses standard ways of optimizing eqs.˜3.49 and 3.50, which could be with a likelihood ratio gradient estimator [Williams, 1992] or via reparameterization.
@@ 3.6.6 Guided policy search (GPS)
The GPS family of methods [Levine and Koltun, 2013, Levine and Abbeel, 2014, Levine et al., 2016, Montgomery and Levine, 2016] fits an approximate dynamics model to data and then amortizes the solutions of a controller that solves the MPC problem in eq.˜3.38 within a local region of the data to improve the amortization model. Given samples of $\pi^{\star}$ from a controller that solves eq.˜3.40, GPS methods typically optimize the KL divergence between the controller and these samples with
@@ 3.6.7 POPLIN by Wang and Ba [2020]
Control. Next they explore ways of using the learned policy $\pi_{\theta^{\star}}$ to solve the model predictive control optimization problem in eq.˜3.38. The POPLIN-A variant (“A” stands for “action”) uses $\pi_{\theta^{\star}}$ to predict an initial control sequence $\hat{u}^{0}_{1:H}$ that is then passed as an input to a controller based on the cross-entropy method over the action space that uses a learned model on the trajectories. The POPLIN-P variant (“P” stands for “parameter”) suggests that the parameter space of the fully-amortized model has learned useful information about the structure of the optimal action sequences. As an alternative to solving the MPC problem in eq.˜3.38 over the action space, POPLIN-P proposes to use CEM to find a perturbation $\omega$ to the optimal parameters $\theta^{\star}$ that maximizes the value from a state $x$ with
@@
Thus the action produced by $\pi_{\theta^{\star}+\omega^{\star}}(x)$ is a solution the control problem in eq.˜3.38 obtained by adapting the policy’s parameters to the state $x$.
@@
The initial phase is fully-amortized behavioral cloning. The second fine-tuning phase can be seen as semi-amortization that learns only the initialization $\theta$ and finds $\omega$ with CEM, with a regression-based loss ${\mathcal{L}}_{\rm reg}$ that only has knowledge of the initial model and does not include the adaptation.
@@ 3.6.8 The differentiable cross-entropy method by Amos and Yarats [2020]
Amos and Yarats [2020] considers a semi-amortized learning setting that learns a latent domain for control, which can be seen as a similarly-motivated alternative to the parameter-space control done in POPLIN. The key piece of latent control is to learn a decoder $\varphi_{\theta}:{\mathcal{Z}}\rightarrow{\mathcal{U}}^{H}$ that maps from a low-dimensional latent space $\mathcal{Z}$} to the $H$-step control space that solves eq.˜3.38. Learning a latent space is useful if there are many redundancies and possibly bad local minima on the original control space ${\mathcal{U}}^{H}$ that the latent space can get rid of. Given an initial state $x_{1}$, the optimal latent representation can be obtained by solving the control optimization problem over $\mathcal{Z}$} with
@@
Solving eq.˜3.56 with DCEM enables the optimal solution $\hat{z}_{\theta}(x)$ to be differentiated with respect to the parameters $\theta$ of the decoder $\varphi_{\theta}$. The final predicted control sequence can be obtained with the decoder $\hat{u}_{1:H}(x;\theta)\vcentcolon=\varphi_{\theta}(\hat{z}_{\theta}(x))$ and the decoder can be learned by regressing onto ground-truth control sequences $u^{\star}(x)$ with
@@
Figure˜3.3 visualizes an example on the cartpole task where this is able to learn a latent space that captures the cyclic and smoothness structure of the optimal control sequence space.
@@ 3.6.10 Learning the value function
The $Q$-value function eq.˜3.36 for finding an MDP policy is often unknown and needs to be estimated from data along with the policy $\pi$ that is amortizing it, e.g. in eq.˜3.37. Actor-critic methods are a common reinforcement learning approach that jointly learn a policy $\pi$ (the actor) and $Q$-value estimate (the critic) [Konda and Tsitsiklis, 1999, Sutton and Barto, 2018]. The policy amortizes the current $Q$ estimate, e.g. by using an approach previously discussed in this section, and the $Q$ function is fit to data sampled from the system. One way of learning the $Q$ function is to replace the value estimate $V^{\pi}$ in eq.˜3.36 with the $Q$-value estimate to yield the relationship
@@
which is referred to as the Bellman equation [Bellman, 1966, Sutton and Barto, 2018]. Eq.˜3.61 is an equality that should hold over all states and actions in the system and a value estimate can be parameterized and learned to satisfy the relationship. While the best way of learning the $Q$ estimate is an open research topic [Watkins and Dayan, 1992, Baird, 1995, Ernst et al., 2005, Maillard et al., 2010, Scherrer, 2010, Geist et al., 2017, Le et al., 2019, Fujimoto et al., 2022], a common way is to optimize residual of eq.˜3.61 with
@@
VAE on MNIST (4.1.1) $784$ ($28\cdot 28$, MNIST digits) 2020 (parameterizing a 10D Gaussian)
@@
Model-based control (4.1.3) $45$ (humanoid states) $51$ ($17\cdot 3$, short action sequence)
Sphere (4.2) $16$ ($c$-convex function parameterizations) 33 (sphere)
Turning now to the implementation details, this section looks at how to develop and analyze amortization software. The standard and easiest approach in most settings is to use automatic differentiation software such as Maclaurin et al. [2015a], Al-Rfou et al. [2016], Abadi et al. [2016], Bezanson et al. [2017], Agrawal et al. [2019b], Paszke et al. [2019], Bradbury et al. [2020] to parameterize and learn the amortization model. There are many open source implementations and re-implementations of the methods in chapter˜3 that provide a concrete starting point to start building on them. This section looks closer at three specific implementations: section˜4.1 evaluates the amortization components behind existing implementations of variational autoencoders section˜3.1 and control section˜3.6 and section˜4.2 implements and trains an amortization model to optimize functions defined on a sphere. Table˜4.1 summarizes the concrete dimensions of the amortization problems considered here and section˜4.3 concludes with other useful software references. The source code behind this section is available at https://github.com/facebookresearch/amortized-optimization-tutorial.
@@ 4.1 Amortization in the wild: a deeper look
This section shows code examples of how existing implementations using amortized optimization define and optimize their models for variational autoencoders (section˜4.1.1) and control and policy learning (sections˜4.1.2 and 4.1.3). The amortization component in these systems is often a part of a larger system to achieve a larger task: VAEs also reconstruct the source data after amortizing the ELBO computation in eq.˜3.5 and policy learning methods also estimate the $Q$-value function in section˜3.6.10. This section scopes to the amortization components to show how they are implemented. I have also added evaluation code to the pre-trained amortization models from existing repositories and show that the amortized approximation often obtains a solution up to 25000 times faster than solving the optimization problems from scratch on an NVIDIA Quadro GP100 GPU.
@@ 4.1.1 The variational autoencoder (VAE)
This section looks at the code behind standard VAE [Kingma and Welling, 2014] that follows the amortized optimization setup described in section˜3.1.1. While there are many implementations for training and reproducing a VAE, this section will use the implementation at https://github.com/YannDubs/disentangling-vae, which builds on the code behind Dupont [2018] at https://github.com/Schlumberger/joint-vae. While the repository is focused on disentangled representations and extensions of the original VAE formulation, this section only highlights the parts corresponding to the original VAE formulation. The code uses standard PyTorch in a minimal way that allow us to easily look at the amortization components.
Training the VAE. Figure˜4.2 paraphrases the relevant snippets of code to implement the main amortization problem in eq.˜3.4 for image data where the likelihood is given by a Bernoulli. Figure˜4.2(a) defines an encoder $\hat{\lambda}_{\theta}$, to predicts a solution to the ELBO implemented in fig.˜4.2(b), which is optimized in a loop over the training data (images) in fig.˜4.2(c). The README in the repository contains instructions for running the training from scratch. The repository contains the binary of a model trained on the MNIST dataset [LeCun, 1998], which the next portion evaluates.
@@
Figure 4.1: Runtime comparison between Adam and an amortized encoder $\hat{\lambda}_{\theta}$ to solve eq.˜3.3 for a VAE on MNIST. This uses a batch of 1024 samples and was run on an unloaded NVIDIA Quadro GP100 GPU. The values are normalized so that $\lambda(x)=0$ takes a value of -1 and the optimal $\lambda^{\star}$ takes a value of 0. The amortized policy is approximately 25000 times faster than solving the problem from scratch.
1class Encoder($n$.Module): # From disvae.models.encoders
@@
(b) Definition of the $\operatorname{ELBO}$ in eq.˜3.1
@@
Evaluating the VAE. This section looks at how well the amortized encoder $\hat{\lambda}$ approximates the optimal encoder $\lambda^{\star}$ given by explicitly solving eq.˜3.3, which is referred to the amortization gap [Cremer et al., 2018]. eq.˜3.3 can be solved with a gradient-based optimizer such as SGD or Adam [Kingma and Ba, 2015]. Figure˜4.4 shows the key parts of the PyTorch code for making this comparison, which can be run on the pre-trained MNIST VAE with code/evaluate_amortization_speed_function_vae.py.
Figure˜4.1 shows that the amortized prediction from the VAE’s encoder predicts the solution to the ELBO 25000 times faster (!) than running 2k iterations of Adam on a batch of 1024 samples. This is significant as every training iteration of the VAE requires solving eq.˜3.3, and a large model may need millions of training iterations to converge. Amortizing the solution makes the difference between the training code running in a few hours instead of a few months if the problem was solved from scratch to the same level of optimality. Knowing only the ELBO values is not sufficient to gauge the quality of approximate variational distributions. To help understand the quality of the approximate solutions, fig.˜4.3 plots out the decoded samples alongside the original data.
@@
Figure 4.4: Evaluation code for comparing the amortized prediction $\hat{y}$ to the true solution $y^{\star}$ solving eq.˜1.1 with a gradient-based optimizer. The full instrumented version of this code is available in the repository associated with this tutorial at code/evaluate_amortization_speed_function.py.
@@ 4.1.2 Control with a model-free value estimate
Figure 4.5: Runtime comparison between Adam and a learned policy $\pi_{\theta}$ to solve eq.˜3.37 on the humanoid MDP. This was evaluated as a batch on an expert trajectory with 1000 states and was run on an unloaded NVIDIA Quadro GP100 GPU. The values are normalized so that $\pi(x)=0$ takes a value of -1 and the optimal $\pi^{\star}$ takes a value of 0. The amortized policy is approximately 1000 times faster than solving the problem from scratch.
@@
This section dives into the training and evaluation code for learning a deterministic model-free policy $\pi_{\theta}:{\mathcal{X}}\rightarrow{\mathcal{Y}}$ to amortize a model-free value estimate $Q$ for controlling the humanoid MDP from Brockman et al. [2016] visualized in fig.˜1.2. This MDP has $|{\mathcal{X}}|=45$ states (angular positions and velocities describing the state of the system) and $|{\mathcal{Y}}|=17$ actions (torques to apply to the joints). A model-free policy $\pi$ maps the state to the optimal actions that maximize the value on the system. Given a known action-conditional value estimate $Q(x,u)$, the optimal policy $\pi^{\star}$ solves the optimization problem in eq.˜3.37 that the learned policy $\pi$ tries to match, e.g. using policy gradient in eq.˜3.47.
@@
The codebase behind Amos et al. [2021] at https://github.com/facebookresearch/svg contains trained model-free policy and value estimates on the humanoid in addition to model-based components the next section will use. The full training code there involves parameterizing a stochastic policy and estimating many additional model-based components, but the basic training loop for amortizing a deterministic policy from the solution there can be distilled into a form similar to fig.˜4.2.
@@
This section mostly focuses on evaluating the performance of the trained model-free policy in comparison to maximizing the model-free value estimate eq.˜3.37 from scratch for every state encountered. An exhaustive evaluation of a solver for eq.˜3.37 would need to ensure that the solution is not overly adapted to a bad part of the $Q$ estimate space — because $Q$ is also a neural network susceptible to adversarial examples, it is very likely that directly optimizing eq.˜3.37 may result in a deceptively good policy when looking at the $Q$ estimate that does not work well on the real system. For simplicity, this section ignores these issues and normalizes the values to [1,0][-1,0] where 1-1 will correspond to the value from taking a zero action and 0 will correspond to the value from taking the expert’s action. (This is valid in this example because the zero action and expert action never coincide.)
@@
Figure˜4.5 shows that the amortized policy is approximately 1000 times faster than solving the problem from scratch. The $Q$ values presented there are normalized and clamped so that the expert policy has a value of zero and the zero action has a value of -1. This example can be run with code/evaluate_amortization_speed_function_control.py, which shares the evaluation code also used for the VAE in fig.˜4.4.